The perimeter of a certain sector of a circle is equal to half circumference if circle. Express the angle of the sector

Grade:11

1 Answers

Nikitha
6 years ago
Let -radius of circle be r -Angle of sector be x -Angle subtend s on the circumferencePerimeter of the sector will be equal to s+2rHalf of circumference is equal to 2πr/2=πr S+2r=πr (given)Rx+2r=πr ( as s=Rx)We get x=π-2 as the relation